# On the Effects of Randomness on Stability of Learning with Limited Labelled Data: A Systematic Literature Review

### v1

@@ Abstract.
Learning with limited labelled data, such as few-shot learning, meta-learning or transfer learning, aims to effectively train a model using only small amount of labelled samples. However, these approaches were observed to be excessively sensitive to the effects of uncontrolled randomness caused by non-determinism in the training process. The randomness negatively affects the stability of the models, leading to large variance in results across training runs. When such instability is disregarded, it can unintentionally, but unfortunately also intentionally, create an imaginary perception of research progress. Recently, this area started to attract a research attention and the number of relevant studies is continuously growing. In this survey, we provide a comprehensive overview of 134 papers addressing the effects of randomness on the stability of learning with limited labelled data. We distinguish between four main tasks addressed in the papers (investigate/evaluate; determine; mitigate; benchmark/compare/report randomness effects), providing findings for each one. Furthermore, we identify and discuss seven challenges and open problems together with possible directions to facilitate further research. The ultimate goal of this survey is to emphasise the importance of this growing research area, which so far has not received appropriate level of attention.
randomness, stability, meta-learning, transfer learning, language model fine-tuning, in-context learning, literature survey
ccs: General and reference Surveys and overviewsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Learning paradigms
@@ 1. Introduction
The approaches for learning with limited labelled data are designed to achieve high performance of the machine learning models even with few labels available (Brown et al., 2020; Wang et al., 2020, 2019c; Song et al., 2023). Under the term learning with limited labelled data we understand any approach that is designed to work with the lack of labels, without any constraint on how the labelled samples are distributed, i.e., whether all the samples are from a single task or are distributed across different tasks. Although similar to the notion of few-shot learning, it represents a broader scope encompassing a larger number of possible approaches (some of which are incorrectly categorised as few-shot learning in current practice).
@@
To deal with the limited labels, the approaches utilise additional information from different sources (Brown et al., 2020; Wang et al., 2020, 2019c). A popular option is to transfer knowledge from similar tasks and datasets using transfer learning. For example, in the NLP domain it is common to use language model fine-tuning, where a language model pre-trained on a large set of unlabelled data (from unrelated domain) is fine-tuned for the specific downstream task or problem statement using only few labelled samples available for it (Xia et al., 2020; Rogers et al., 2020; Mosbach et al., 2021; Dodge et al., 2020). Another emerging technique is the in-context learning, where a large pre-trained language model is “prompted“ to predict a label for a single test sample by presenting it with a concatenation of few labelled samples and the test sample, without requiring any parameter update (Brown et al., 2020; Liu et al., 2023). Finally, meta-learning is often used, where the model is explicitly trained to quickly adapt to a new task with only a handful of examples by learning how best to learn across large number of related tasks with few labelled samples each (Agarwal et al., 2021; Hospedales et al., 2021; Kadam and Vaidya, 2018).
@@
However, a significant problem observed for the approaches for learning with limited labelled data is their sensitivity to the effects of uncontrolled randomness, which negatively affects their stability. Under stability, we understand a property of a learning algorithm or a model that indicates what influence the small scale and random changes (or perturbations) in the input data and parameters have on its outputs (Liu et al., 2017). Such random changes are introduced by different randomness factors that represent the non-deterministic decisions in the training process, such as random initialisation of model parameters, or data shuffling (Pham et al., 2021; Gundersen et al., 2022). These randomness factors represents the main point around which the effects of randomness on the stability are addressed.
@@
The uncontrolled effects of randomness can have massive impact on the stability of the utilised machine learning approaches. Running the training multiple times on the same dataset, with the same setup and hyperparameters may lead to large deviation in the final performance (Mosbach et al., 2021; Dodge et al., 2020; Lu et al., 2022; Agarwal et al., 2021). Repeating BERT fine-tuning multiple times and cherry-picking the best result, it is possible to outperform more recent BERT variants (Dodge et al., 2020). Changing only order of data in in-context learning can lead the model from state-of-the-art predictions to random guesses (Lu et al., 2022). Or choosing different set of adaptation data in meta-learning can lead to difference between minimum and maximum performance being up to 90% (Agarwal et al., 2021). Even though the effects of randomness are present in all cases of machine learning, their impact on the stability is especially significant in low data regimes.
@@
If the uncontrolled randomness is not appropriately addressed in such low data regimes, it can have a significant non-negligible negative consequences. In research works and benchmarks, it may prohibit an objective performance comparison of newly designed methods to the state-of-the-art baselines (especially when the difference between their performance is similar to the deviation caused by the randomness). In such case, a method can be incorrectly denoted as state-of-the-art only based on a more favourable random chance (Reimers and Gurevych, 2017) (as illustrated in Figure 0(b)). The uncontrolled randomness can unintentionally, but unfortunately also intentionally (by cherry-picking), create an imaginary perception of research progress.
The effects of randomness are addressed in various depths - from a pure recognition of randomness effects, through a deeper investigation up to a full mitigation of such effects (illustrated in the Figure 1). Recent few-shot learning surveys recognise the effects of randomness as an open problem (Wang et al., 2020; Tian et al., 2022; Dong et al., 2022), or even dedicate a part of the survey to a basic overview of these effects (Song et al., 2023). Other papers simply recognise the impact of randomness on the overall results and explicitly repeat the training and evaluation multiple times to obtain a more representative, albeit aggregated value to address the randomness. However, no further analysis regarding the effects of the randomness is provided. Finally, only a small fraction of papers provide more detailed analysis of the effects of randomness to improve the stability and reduce the significant deviation in results. These works conduct a series of experiments that are designed to investigate the effects, estimate the distribution of results and determine the significance and the impact of the different randomness factors on the stability. The effects of the randomness are analysed in more detail, in order to determine the real origin of randomness in the training process (e.g., under-specification, which causes slightly different parameter initialisation to converge to different minima). Based on the findings from the analysis of the experiments, various methods are proposed to mitigate the effects of randomness, improving the stability and reducing the deviation to achieve more stable results.
@@
Figure 1. The effects of randomness are addressed in various depths in the papers. First, the effects are simply recognised, resulting in reporting of a single aggregated value from multiple repeated training and evaluation runs (illustrated as “Recognise”). This leads to biased comparison between approaches, as one approach may show better results only due to random chance and unintentional cherry-picking (illustrated in Figure 0(b)). Analysing the effects of randomness, estimating the distribution of results from multiple runs and comparing the distributions can prevent this problem (illustrated as “Describe”). As the distribution is not affected in any way, only estimated, the deviation is not reduced. The effects of randomness are properly addressed when they are mitigated, reducing the deviation in the distribution of results (illustrated as “Mitigate”). However, an effective mitigation requires understanding of the effects, their importance, and the origin of randomness provided by the analysis of the effects of randomness (provided by the “Describe” task).
@@
In this paper, we provide a comprehensive survey of 134 papers that address the effects of randomness, either as simply recognising the effects of randomness on the stability (N=84), or investigating and describing the characteristics of these effects in more detail and/or exploring the appropriate means of their mitigation (N=50). We provide a more in-depth analysis of the 50 papers that address the effects of randomness in more detail. We provide aggregated findings from this analysis and use them to identify seven open problems related to the study of the effects of randomness, while providing future directions on how to address them.
@@
This is the first survey of its kind that provides a comprehensive and systematic literature review specifically focused on the impact of the effects of randomness on stability across various approaches for dealing with limited labelled data. As opposed to previous papers that provide an overview of possible randomness factors and investigate these effects (Pham et al., 2021; Gundersen et al., 2022), we focus on all tasks for addressing the effects of randomness, not only their investigation, but also determining their origin, mitigation and benchmarking in presence of these effects of randomness.
@@
The purpose of this survey is to emphasise the importance of this research area, which so far has not received appropriate level of attention from researchers and practitioners. At first, it should serve existing or new-coming researchers who are explicitly tackling with the topic of the randomness and its effects on the stability of learning with limited labelled data to support their research (by providing the overview of the state-of-the-art, identification of open problems and future direction). Secondly, its purpose is also to inform researchers and practitioners utilising the learning with limited labelled data about the consequences of unaddressed randomness and how to effectively prevent them. To better achieve these purposes, we summarised all the identified papers along with their categorisation as in the supplementary material of this survey (Section A).
@@
The rest of the paper is structured as follows. In Section 2, we describe the scope of the survey, difference to existing surveys and the methodology used for identifying relevant papers. Taxonomy applied to analyse and categorise the identified papers is described in Section 3. Afterwards, we describe the different tasks for addressing the effects of randomness, categorised based on our taxonomy, in Sections 4, 5, 6 and 7, with a summary of the findings at the end of each such section. The identified challenges and open problems are described in Section 8. Finally, the survey is concluded in Section 9 with summary of our contributions and findings.
@@ 2.1. Scope of the Survey
In this survey, we focus on a specific part of stability. We focus on papers that explore the effects of small perturbations to multiple inputs of the training process on the final performance after the training process is done (as opposed to the effects of the perturbation of a single sample on the loss or gradient of training as defined by uniform stability (CHEN et al., 2020)). A similar, but at the same time quite distinct property is the robustness of the model. It deals with a more large scale or systematic changes in the input, such as distribution shift or presence of adversarial examples. As the adversarial robustness is already extensively studied in the context of learning with limited labelled (e.g., in meta-learning (Goldblum et al., 2020; Yin et al., 2018; Xu et al., 2021a; Wang et al., 2021b)) and the effects are often systematic, we consider it to be out of scope for this survey. Finally, we view the stability to be a subset of the reproducibility of the model. Besides randomness in the training process, the reproducibility is impacted by other factors such as selective reporting, biases, data and source code availability, and many others as specified in (Gundersen et al., 2022). These additional factors are out of scope for this survey.
In case of machine learning approaches, we mainly focus on the approaches that utilise related tasks and datasets as additional information. Not requiring large set of labelled data, these approaches are usable across broader set of domains, making them also more popular. Due to this popularity, they also cover almost the whole research area of addressing the effects of randomness on the stability, as there are almost no other works focusing on approaches that utilise information from the unlabelled sets of data. Our focus is on three prevalent groups of approaches: 1) meta-learning; 2) language model fine-tuning; and 3) in-context learning. Even though we do not explicitly focus on a specific modality, the popularity of the approaches impact the scope of this survey. As the majority of the papers that focus on stability of transfer learning work with language models, we define the scope in this survey using the language model fine-tuning and in-context learning approaches for better readability. We provide only the general problem statement for each group of approaches (in Section 3.3), while providing references to more comprehensive overview.
@@ 2.2. Existing Surveys
To the best of our knowledge, there are no existing surveys focused on addressing the effects of randomness on stability. However, there are surveys on related topics that partially touch on the topics of stability. First of all, some papers provides an overview of possible sources of randomness, although with focus on replicability (Gundersen et al., 2022). In some surveys on the topic of learning with limited labelled data and/or few-shot learning, the stability and the effects of randomness are mentioned as a problem that should be addressed, with varying level of detail provided (Li et al., 2021; Tian et al., 2022; Dong et al., 2022; Song et al., 2023).
At the same time, the study of effects of randomness on the stability is becoming a research topic addressed in an increasing number of research works focused on dealing with the lack of labels, mainly focused on the use of pre-trained models (Dodge et al., 2020; Mosbach et al., 2021; Lu et al., 2022). Even first benchmarks consider the stability of the models to be paramount and include it as a part of the final metric (Efrat et al., 2022).
To fill this gap, in this survey we aim to provide the first comprehensive overview of the effects of randomness on the stability of learning with limited labelled data and describe it from perspectives that are not easily visible from the standard analysis of related works.
@@ 2.3. Survey Methodology
To identify the set of relevant papers for our comprehensive survey, we take inspiration from the PRISMA methodology (Moher et al., 2009; Page et al., 2021), performing following steps (as illustrated in Figure 2):
@@
    Identification of relevant papers. A keyword search query is defined based on the related terms for our defined scope identified in different papers. Multiple digital libraries are searched using the defined query to identify set of potentially relevant papers.
@@
    Reference analysis. In order to discover additional potentially relevant papers, the papers cited in the core papers; as well as papers citing the core papers, are examined and their relevance is determined based on their title and abstract. The papers deemed potentially relevant are an additional input into the second step of this process.
@@
    Analysis and categorisation. The identified core papers are analysed and categorised based on the taxonomy (as defined in Section 3) to provide the comprehensive overview and to identify challenges and open problems.
As a result, 134 papers are identified — out of them, 84 papers are classified as recognise papers, and 50 as core papers. The number of identified papers grouped by the publication year is presented in Figure 3. The categorisation of the core papers is showcased in Table 1. Moreover, the full list of all identified papers (including recognise papers), with their categorisation and additional metadata is available in the supplementary material (Section A). For a more detailed implementation of the survey methodology, see the supplementary material (Section C).
@@
Figure 3. Number of papers dealing with randomness grouped by the year. Figure 2(a) shows only the core papers that focus on addressing the effects of randomness in more detail, while Figure 2(b) also includes the papers that only recognise the problem. The problem started to attract attention only in recent years.
Table 1. A showcase of the core papers and their categorisation based on the proposed txonomy. The full categorisation, along with the recognise papers is available in the supplementary material (Section A).
References Area Factor Performed Tasks Modality Interactions Out-Of-Distribution

### v2

# A Survey on Stability of Learning with Limited Labelled Data and its Sensitivity to the Effects of Randomness

@@ Abstract.
Learning with limited labelled data, such as prompting, in-context learning, fine-tuning, meta-learning or few-shot learning, aims to effectively train a model using only a small amount of labelled samples. However, these approaches have been observed to be excessively sensitive to the effects of uncontrolled randomness caused by non-determinism in the training process. The randomness negatively affects the stability of the models, leading to large variances in results across training runs. When such sensitivity is disregarded, it can unintentionally, but unfortunately also intentionally, create an imaginary perception of research progress. Recently, this area started to attract research attention and the number of relevant studies is continuously growing. In this survey, we provide a comprehensive overview of 415 papers addressing the effects of randomness on the stability of learning with limited labelled data. We distinguish between four main tasks addressed in the papers (investigate/evaluate; determine; mitigate; benchmark/compare/report randomness effects), providing findings for each one. Furthermore, we identify and discuss seven challenges and open problems together with possible directions to facilitate further research. The ultimate goal of this survey is to emphasise the importance of this growing research area, which so far has not received an appropriate level of attention, and reveal impactful directions for future research.
randomness, stability, sensitivity, meta-learning, large language models, fine-tuning, prompting, in-context learning, instruction-tuning, prompt-based learning, PEFT, literature survey
copyright: acmlicensedjournal: CSURjournalyear: 2024journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3691339ccs: General and reference Surveys and overviewsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Learning paradigms
@@ 1. Introduction
The approaches for learning with limited labelled data are designed to achieve high performance in machine learning models even with few labels available (Liu et al., 2023a; Song et al., 2023). Under the term learning with limited labelled data, we understand any approach that is designed to work with a lack of labels, without any constraint on how the labelled samples are distributed, i.e., whether all the samples are from a single task or are distributed across different tasks. Although similar to the notion of few-shot learning, it represents a broader scope encompassing a larger number of possible approaches (some of which are incorrectly categorised as few-shot learning in current practice).
@@
To deal with the limited labels, these approaches utilise additional information from different sources (Brown et al., 2020; Song et al., 2023), such as transferring knowledge from similar tasks and datasets. In the NLP domain, prompting and in-context learning (also called few-shot prompting) have recently emerged. In these techniques, a large pre-trained language model is “prompted“ to predict a label for a single test sample by presenting it with task instructions and the test sample (and concatenation of a few labelled samples when in-context learning is used), without requiring any parameter update (Liu et al., 2023a). In addition, it is common to use fine-tuning, where the parameters, or their subset using parameter-efficient fine-tuning (PEFT) methods, of the pre-trained large language model are updated to optimise the model for the specific downstream task using only a few labelled samples (Mosbach et al., 2021; Dodge et al., 2020; Chen et al., 2022a). Finally, meta-learning can be used, where the model is explicitly trained to quickly adapt to a new task with only a handful of examples by learning how best to learn across a large number of related tasks with few labelled samples each (Agarwal et al., 2021; Hospedales et al., 2021).
@@
However, a significant problem observed for these approaches is their sensitivity to the effects of uncontrolled randomness, which negatively affects their stability. Under stability and its opposite term sensitivity, we understand a property of a learning algorithm or a model that indicates what influence the small-scale and random changes (or perturbations) in the input data and parameters have on its outputs. Such random changes are introduced by different randomness factors that represent the non-deterministic decisions in the training process, such as random initialisation of model parameters or data shuffling (Pham et al., 2021; Gundersen et al., 2022). These randomness factors represent the main point around which the effects of randomness are addressed.
@@
The uncontrolled effects of randomness can have a massive impact on the stability of the utilised machine learning approaches. Running the training multiple times on the same dataset, with the same setup and hyperparameters, may lead to large deviations in the final performance (Mosbach et al., 2021; Dodge et al., 2020; Lu et al., 2022; Agarwal et al., 2021). Changing only the order of samples or answers in multi-choice question answering when using in-context learning can lead the model from state-of-the-art predictions to random guesses (Lu et al., 2022; Zong et al., 2023; Wei et al., 2024). Choosing a different set of adaptation data in meta-learning can lead to a difference between minimum and maximum performance being up to 90% (Agarwal et al., 2021). Even though the effects of randomness are present in all cases of machine learning, their impact on stability is especially significant in low-data regimes.
@@
If the uncontrolled randomness is not appropriately addressed in such low-data regimes, it can have significant, non-negligible negative consequences. In comparisons and benchmarks, changing only the random seed may lead to completely different model rankings (Madaan et al., 2024; Alzahrani et al., 2024). It may prohibit an objective performance comparison of newly designed methods to the state-of-the-art baselines, as it makes it difficult to conclude if a specific change made a meaningful difference or if the model just ”got lucky” (especially when the difference between their performance is similar to the deviation caused by the randomness). In such cases, a method can be incorrectly denoted as state-of-the-art only based on a more favourable random chance (Reimers and Gurevych, 2017) (as illustrated in Figure 1(b)). The uncontrolled randomness can unintentionally, but unfortunately also intentionally (by cherry-picking), create an imaginary perception of research progress. As a result, researchers even suggest that we should be very vigilant when relying on models and approaches that are significantly affected by sensitivity (e.g., large language models) (Zong et al., 2023). Finally, randomness has been identified as a significant obstacle that negatively affects reproducibility (Albertoni et al., 2023; Chen et al., 2022b).
The effects of randomness are addressed in various depths - from a pure recognition of the effects, through a deeper investigation and determining the origin of the effects, up to their partial or full mitigation (illustrated in the Figure 1). First, the effects are simply recognised, resulting in the reporting of a more representative, albeit single, aggregated value from multiple training and evaluation runs (denoted as “Recognise”). As no further analysis is provided, this still leads to a biased comparison between approaches, as one approach may show better results only due to random chance and unintentional cherry-picking (illustrated in Figure 1(b)). Another group of works perform analyses of the effects of randomness, estimate the distribution of results from multiple runs and compare the distributions (denoted as “Investigate/Determine”). The effects of the randomness are often analysed in more detail in order to determine the real origin of randomness in the training process (e.g., under-specification, which causes slightly different parameter initialisation to converge to different minima). As the distribution is only estimated, the deviation is still not reduced. The effects of randomness are fully addressed when they are mitigated, reducing the deviation in the distribution of results (denoted as “Mitigate”). However, effective mitigation requires an understanding of the effects, their importance, and the origin of randomness provided by the analysis of the effects of randomness (provided by the “Investigate/Determine” task).
@@
Figure 1. (a) The effects of randomness can be addressed in various depths in the papers. (b) If not taken into consideration, randomness can introduce bias into comparison results, causing one approach to show better results only due to random chance and unintentional cherry-picking.
@@
In this paper, we conduct a comprehensive survey of 415 papers that address the effects of randomness, either by simply recognising the effects of randomness on stability (N=254), or by investigating/determining the characteristics of these effects in more detail and/or exploring the appropriate means of their mitigation (N=161) [^1] The digital appendix with a full list of papers is available at https://kinit.sk/public/acm-csur-sensitivity-survey.html. We provide a more in-depth analysis of the 161 papers that address the effects of randomness in more detail. Finally, we aggregate findings from this analysis and use them to identify seven challenges and open problems related to the study of the effects of randomness, while providing future directions on how to address them. For the purpose of this survey, we follow the PRISMA methodology (Moher et al., 2009; Page et al., 2021) to systematically identify the papers in the defined scope.
@@
This is the first survey of its kind that provides a comprehensive and systematic literature review specifically focused on the impact of the effects of randomness on stability across various approaches for dealing with limited labelled data. As opposed to previous papers that provide an overview of possible randomness factors and investigate these effects (Pham et al., 2021; Gundersen et al., 2022; Song et al., 2023), provide an overview for a single task (e.g., mitigating the effects) (Xu et al., 2024), or only recognise the problem randomness as an open problem (Song et al., 2023; Liu et al., 2023a), we focus on all tasks for addressing the effects of randomness, including investigation, but also determining their origin, mitigation, and benchmarking in the presence of these effects of randomness.
@@
The purpose of this survey is to emphasise the importance of this research area, which so far has not received the appropriate level of attention from researchers and practitioners. First, it should serve existing or new researchers who are explicitly tackling the topic of randomness and its effects on the stability of learning with limited labelled data to support their research (by providing an overview of the state-of-the-art, identification of open problems and future directions). Secondly, its purpose is to inform researchers and practitioners utilising the learning with limited labelled data about the consequences of unaddressed randomness and how to effectively prevent them. To better achieve these purposes, we summarised all the identified papers along with their categorisation in the digital appendix of this survey (which is further described in Section A of the supplementary material).
@@
The rest of the paper is structured as follows. In Section 2, we describe the scope of the survey, the difference to existing surveys and the methodology used for identifying relevant papers. Taxonomy applied to analyse and categorise the identified papers is described in Section 3. Afterwards, we describe the different tasks for addressing the effects of randomness, categorised based on our taxonomy, in Sections 456 and 7, with a summary of the findings at the end of each such section. The identified challenges and open problems are described in Section 8. Finally, the survey is concluded in Section 9 with a summary of our contributions and findings.
@@ 2.1. Scope of the Survey
In this survey, we focus on a specific part of stability and sensitivity of machine learning approaches. We focus on papers that explore the effects of small perturbations to multiple inputs of the training process on the final performance after the training process is done. A similar, but at the same time quite distinct property is the robustness of the model. It deals with a more large scale or systematic changes in the input, such as distribution shift or the presence of adversarial examples. As the adversarial robustness is already extensively studied in the context of learning with limited labelled and the effects are often systematic, we consider it to be out of scope for this survey. Finally, we view the stability to be a subset of the reproducibility of the model. Besides randomness in the training process, reproducibility is impacted by other factors such as selective reporting, biases, data and source code availability, and many others as specified in (Gundersen et al., 2022). These additional factors are out of scope for this survey.
Out of a wide spectrum of machine learning approaches, we specifically focus on learning with limited labelled data, which is designed to train or prompt a model with only a limited number of labelled samples (typically 0-50 per class). Not requiring large sets of labelled data, these approaches are usable across a broader set of domains, making them also more popular. Due to this popularity, they also cover almost the whole research area of addressing the effects of randomness on stability, as there are almost no other works focusing on approaches that utilise information from unlabelled sets of data. Our focus is on five prevalent groups of approaches: 1) meta-learning; 2) language model (LM) fine-tuning; 3) prompting/in-context learning; 4) prompt-based learning; and 5) parameter-efficient fine-tuning (FT). Even though we do not explicitly focus on a specific modality, the popularity of the approaches impacts the scope of this survey. The most prominent modality is text, followed by images, as the majority of the papers focus on language model fine-tuning and in-context learning. Finally, even though we focus on the setting of limited data, we also include works that are relevant for this survey that utilise large sets of data. However, due to the search query used, such papers are not systematically covered and may be incomplete (even though we specifically designed one step in the methodology to catch these papers).
@@ 2.2. Existing Surveys
To the best of our knowledge, there are no existing surveys addressing the effects of randomness on stability across all tasks ranging from recognise to mitigate. However, there are relevant surveys on related topics that partially touch on the topics of stability or sensitivity. First of all, some papers provide an overview of possible sources of randomness, although with a focus on replicability (Gundersen et al., 2022). In some surveys on learning with limited labelled data and/or few-shot learning, the sensitivity to the effects of randomness is already mentioned as a problem that should be addressed (Dong et al., 2022; Song et al., 2023; Liu et al., 2023a; Kaddour et al., 2023), but no additional details are provided. Finally, the most recent relevant survey (Xu et al., 2024) focuses on sample selection strategies for mitigating in-context learning sensitivity to sample choice. While relevant, it provides a significantly narrower perspective than the one given by this comprehensive survey.
To fill this gap, we aim to conduct the first comprehensive overview of the effects of randomness on the stability of learning with limited labelled data and provide novel perspectives that are not easily visible from the standard analysis of related works.
@@ 2.3. Survey Methodology
To systematically identify the set of relevant papers for our comprehensive survey, we follow the PRISMA methodology (Moher et al., 2009; Page et al., 2021), performing the following steps (as illustrated in Figure 2):
@@
    Identification of relevant papers. A keyword search query is defined based on the related terms for our defined scope identified in different papers. Multiple digital libraries are searched using the defined query to identify a set of potentially relevant papers.
@@
    Reference analysis. To discover additional potentially relevant papers (especially those relevant for the survey that do not necessarily deal with limited data), the papers cited in the core papers; as well as papers citing the core papers, are examined and their relevance is determined based on their title and abstract. The papers deemed potentially relevant are an additional input into the second step of this process.
@@
    Analysis and categorisation. The identified core papers are analysed and categorised based on the taxonomy (as defined in Section 3) to provide a comprehensive overview and to identify challenges and open problems.
As a result, 415 papers are identified — out of them, 254 papers are classified as recognise papers, and 161 as core papers. The number of identified papers grouped by the publication year is presented in Figure 3. The categorisation of the selected core papers is showcased in Table 1. Moreover, the full list of all identified papers (including recognise papers), with their categorisation and additional metadata is available in the digital appendix. For a more detailed implementation of the survey methodology, see the supplementary material (Section C).
@@
Figure 3. Number of papers dealing with randomness grouped by the year. Figure 3(a) shows only the core papers that focus on addressing the effects of randomness in more detail, while Figure 3(b) also includes the papers that only recognise the problem. The problem started to attract attention only in recent years. Year 2024 covers papers published until July 2024.
Table 1. A showcase of the selected core papers and their categorisation based on the proposed taxonomy. The full categorisation, along with the recognise papers is available in the digital appendix.
References ML Approach Randomness Factor Performed Tasks Modality Interactions Out-Of-Distribution